# VLMEvalKit: An Open-Source Toolkit for Evaluating Large Multi-Modality Models

### v3

@@ 2 VLMEvalKit: Design & Features
Currently, a multi-modal message primarily incorporates image and text modalities. The format, however, is extensible and can accommodate additional modalities such as audio or point clouds.
@@
Multi-modal Inference. To expedite the multi-modal inference, we support parallelized inference for both commercial APIs, leveraging Python’s multiprocessing, and open-source models, which are distributed in parallel across multiple GPUs. Additionally, we have implemented a robust inference process so that an interrupted inference process can be resumed with minimal costs of repeated calculation or API calls.
@@
HunYuan-Standard-Vision [95] N/A 76.3 83.6 72.9 70.7 66.1 57.7 93.2 861 79.9
Qwen2.5-VL-72B [5] 73.4B 76.2 87.8 71.1 67.9 70.8 58.8 88.2 881 76.7
TeleMM [94] N/A 75.9 79.9 70.8 66.6 75.7 60.6 88.5 891 75.7
InternVL2.5-38B-MPO [12] 38B 75.3 85.4 70.1 63.8 73.6 59.7 87.9 894 72.6
InternVL2.5-78B [12] 78B 75.2 87.5 69.5 70.0 70.6 57.4 89.1 853 71.8
Table 2: The evaluation results of LMMs on general VQA benchmarks. The table displays the top-10 LMMs, including both commercial APIs and open-source LMMs(till 2025.02.28), in the descending order of average score. When calculating the average score, scores of each benchmark are normalized to the range of 0 to 100. When reporting the parameter size, ‘1B’ means $10^{9}$ parameters. Commercial APIs are denoted with blue background. Bold and underline indicates the best and second-best performance in each group.
@@
Gemini-1.5-pro-002 [92] N/A 50 67.9 41 54.8 31.5 50.5 54.4
GLM-4v-Plus-20250111 [3] N/A 49.2 73.5 51.1 40.7 27.5 47.7 54.4
Claude-3.5-sonnet-20241022 [4] N/A 47.9 65.3 35.6 46.3 35.7 44 60.4
Ovis2-34B [68] 34B 47.9 76.1 31.9 50.1 27.5 51.9 49.9
QVQ-72B-Preview [5] 72B 46.9 70.3 34.9 48.2 30.7 39 58.2
Step-1o [87] N/A 45.4 74.7 27.9 42 29.7 45.3 52.8
Ovis2-16B [68] 16B 44.7 73.7 30.1 45.8 26.3 45 47.4
InternVL2.5-78B-MPO [12] 78B 44.3 76.6 36.2 43.7 21.2 37.6 50.8
@@ 3.1 General VQA Benchmarks
As illustrated in the table, after nearly two years of development, the performance gap between commercial APIs and open-source LMMs on general VQA benchmarks has significantly narrowed. Open-source LMMs now demonstrate strong capabilities in general understanding tasks, often matching or even surpassing the performance of commercial APIs. Among the top-10 models, five are commercial APIs, while the other five are open-source models. Although the top two models, Step-1o [87] and SenseNova [83], remain commercial APIs, their performance is remarkably close to that of the third-ranked model, InternVL2.5-78B-MPO [12], with scores of 77.7%, 77.4%, and 77.0% on average, respectively. From the first-ranked model, Step-1o, to the tenth-ranked model, InternVL2.5-78B, the average score decreases by only 2.5% (from 77.7% to 75.2%). Notably, prominent commercial series such as GPT [77], Gemini [92], and Claude [4] fail to rank within the top 10, whereas open-source models like the InternVL and QwenVL series exhibit exceptional performance and competitive strength.
@@ 3.2 Image Reasoning Benchmarks
In multi-modal reasoning tasks, a significant performance gap remains evident between commercial APIs and open-source LMMs. Generally, commercial APIs demonstrate a marked advantage over their open-source counterparts. Notably, the top five performing LMMs are all proprietary APIs. Ovis2-34B [68], the highest-performing open-source LMM on reasoning benchmarks, achieves an average score of only 47.9, highlighting a considerable gap compared to the leading model Doubao-1.5-pro [91](61.6). Doubao-1.5-pro, surpasses the second-ranked model, Gemini-2.0-flash, by a substantial margin of +11.0 on average (61.6 vs. 50.6). Further analysis reveals that most high-performing models achieve accuracy exceeding 70% on MathVista[66]. In contrast, only Doubao-1.5-pro and Gemini-2.0-flash achieve accuracy greater than 40% on DynaMath [133], underscoring the challenges posed by this benchmark. Despite its leading performance in reasoning tasks, Doubao-1.5-pro achieves an average accuracy of just 61.6%, indicating that there is still substantial room for improvement. These findings suggest that the reasoning capabilities of LMMs require further exploration and development.
@@ 4 Discussion
We have released VLMEvalKit, an open-source toolkit designed for the evaluation of large multi-modality models. VLMEvalKit encompasses a comprehensive collection of over 200 LMMs and more than 80 multi-modal benchmarks. The codebase is structured with simplicity in mind, facilitating the easy integration of new LMMs or benchmarks. Our framework is thoughtfully designed to extend its capabilities beyond the image modality, we have recently incorporated the evaluation of a video understanding benchmark, MMBench-Video [19], into VLMEvalKit. Moving forward, the development of VLMEvalKit will focus on expanding the repertoire of LMMs and benchmarks for video and other modalities. We are optimistic that this repository, along with all released resources, will contribute to advancing research in multi-modal learning.
@@ B.1 Contributors with 3+ Major Contributions
  • PhoenixZ810 (Xiangyu Zhao): The contributor helped support WeMath [81], LogicVista [112], MM-AlignBench [132], Video-ChatGPT [72], Chat-UniVI [39], and Llama-VID [52].
  • amitbcp (Amit Agarwal): The contributor helped support MUIRBench [98], Phi-3.5 [1], Idefics3 [42], VILA [56], xGen-MM [114], and MVTamperBench [2].
@@ B.2 Full Contributor List
  • mary-0830: The contributor supports OmChat [131].
@@
  • lihytotoro: The contributor supports MiniCPM-V-2.6 [117].
@@
  • FeipengMa6: The contributor supports WeMM [17].
@@
  • lrlbbzl: The contributor of URSA-8B [69].
  • lyccnb, jscslld: The contributors of CG-Bench [6].
@@
  • wulipc: The contributor of CC-OCR [116].
  • mfarre: The contributor of SmolVLM.
@@
  • CaraJ7: The contributor of MathVerse [125].

### v4

@@ 2 VLMEvalKit: Design & Features
Currently, a multi-modal message primarily incorporates image, video and text modalities. The format, however, is extensible and can accommodate additional modalities such as audio or point clouds.
@@
Multi-modal Inference. To expedite the multi-modal inference, we support parallelized inference for both commercial APIs, leveraging Python’s multiprocessing, and open-source models, which are distributed in parallel across multiple GPUs. Additionally, we have implemented a robust inference process so that an interrupted inference process can be resumed with minimal costs of repeated calculation or API calls. For some popular open-source LMMs(like Qwen Series, InternVL Series), we have supported vllm/LMDeploy to accelerate the inference with full utilization of resources.
@@
Table 2: The evaluation results of LMMs on general VQA benchmarks. The table displays the top-10 LMMs, including both commercial APIs and open-source LMMs(till 2025.06.20), in the descending order of average score. When calculating the average score, scores of each benchmark are normalized to the range of 0 to 100. When reporting the parameter size, ‘1B’ means $10^{9}$ parameters. Commercial APIs are denoted with blue background. Bold and underline indicates the best and second-best performance in each group.
@@
Claude-3.7-sonnet-20250219 [6] N/A 50.4 66.8 41.9 46.7 39.7 49.3 58.2
InternVL3-38B [189] 38B 50.3 76.3 35.9 48.2 34.7 49.3 57.7
@@ 3.1 General VQA Benchmarks
As illustrated in the table, after nearly two years of development, the performance gap between commercial APIs and open-source LMMs on general VQA benchmarks has significantly narrowed. Open-source LMMs now demonstrate strong capabilities in general understanding tasks, often matching or even surpassing the performance of commercial APIs. Among the top-10 models, four are commercial APIs, while the other five are open-source models. Although the top three models, CongRong-v2.0 [24], SenseNova-V6-Pro [108], and Gemini-2.5-Pro [121] remain commercial APIs, their performance is remarkably close to that of the fourth-ranked model, InternVL3-78B [189]. From the first-ranked model, CongRong-v2.0, to the tenth-ranked model, Ovis2-34B, the average score decreases by only 4.2% (from 80.7% to 76.5%). Notably, prominent commercial series such as GPT [99] and Claude [6] fail to rank within the top 10, whereas open-source models like the InternVL and Ovis series exhibit exceptional performance and competitive strength.
Method Param. Frame. Avg. Score MVBench Video-MME(w/o subs) MMBench-Video TempCompass MLVU
InternVL3-78B [189] 78.4 64 72.7 79.2 73.1 1.81 77.0 74.0
InternVL3-38B [189] 38.4 64 71.9 76.0 72.1 1.80 78.5 72.7
Qwen2.5-VL-72B [7] 73.4 64 68.6 71.3 68.6 1.84 77.7 64.1
InternVL2.5-78B [18] 78.4 64 68.1 75.6 72.1 1.98 56.0 70.9
Qwen2-VL-72B [137] 73.4 64 67.7 70.2 67.3 1.70 79.4 65.2
InternVL3-8B [189] 7.94 64 66.8 73.2 66.0 1.69 70.4 68.0
LLaVA-Video-72B-Qwen2 [179] 7.94 64 66.2 63.1 70.5 1.71 70.4 70.0
Aria [66] 25.3 64 66.2 67.9 66.0 1.81 69.6 67.1
Gemini-2.0-flash [121] N/A 64 64.6 59.4 71.0 2.01 57.9 68.0
GPT-4o-20240806 [99] N/A 16 63.1 57.5 67.9 1.87 72.7 54.9
Qwen2.5-Omni-7B [155] 10.7 64 64.5 69.0 64.1 1.65 70.7 63.6
Table 4: The evaluation results of LMMs on video understanding benchmarks. The table displays the top-10 LMMs, including both commercial APIs and open-source LMMs(till 2025.06.25), in the descending order of average score. Due to resource limitations, we only tested a few commercial APIs.
@@ 3.2 Image Reasoning Benchmarks
In multi-modal reasoning tasks, a significant performance gap remains evident between commercial APIs and open-source LMMs. Generally, commercial APIs demonstrate a marked advantage over their open-source counterparts. Notably, the top six performing LMMs are all proprietary APIs. InternVL3-78B [189], the highest-performing open-source LMM on reasoning benchmarks, achieves an average score of only 51, highlighting a considerable gap compared to the leading model Seed1.5-VL [42](73.3). Further analysis reveals that most high-performing models achieve accuracy exceeding 85% on MathVista[87]. In contrast, only Seed1.5-VL and Gemini-2.5-Pro achieve accuracy greater than 50% on DynaMath [190], underscoring the challenges posed by this benchmark.
3.3 Video Benchmarks
Video has become essential in daily life, driving communication, learning, and entertainment. With the rapid rise of video content and progress in LMM capability, video understanding ability is increasingly vital for LMMs. To fully examine the video understanding capability, we have added support for some popular common video understanding benchmarks, including MMBench-Video [31], Video-MME [34], MLVU [188], TempCompass [81], MVBench [68], and other video datasets. Besides, in order to have a glimpse into the overall video understanding performance of LMMs, we also selected some representative video benchmarks and evaluated over forty LMMs, with the results presented on the OpenVLM Video Leaderboard. The selected benchmarks primarily cover various question types (multiple choice, open-ended questions, true/false questions, etc.) and various video lengths, mainly examining the model’s general video understanding ability. The performance metrics of the top-10 models are summarized in Tab. 4.
Due to the limited resources of commercial APIs and the significant costs associated with video understanding evaluations, we focused on testing the video understanding capabilities of open-source LMMs. It’s easy to observe that InternVL and Qwen Series demonstrate remarkable performance in video understanding benchmarks. InternVL3-78B [189] achieves the highest average score of 72.7, and Qwen2.5 lags slightly behind them, mainly revealed in Video-MME [34] and MLVU [188]. With a 64-frame input window, Gemini-2.0-flash [121] still lags behind the state of the art on several tasks, possibly because it refuses to answer some questions. GPT-4o [99] achieves 63.1 with only 16 frames, showing that commercial APIs can still compete in video understanding even under a short context.
@@ 4 Discussion
We have released VLMEvalKit, an open-source toolkit designed for the evaluation of large multi-modality models. VLMEvalKit encompasses a comprehensive collection of over 300 LMMs and more than 100 multi-modal benchmarks. The codebase is structured with simplicity in mind, facilitating the easy integration of new LMMs or benchmarks. Our framework is thoughtfully designed to extend its capabilities beyond the vision modality and incorporate other modalities like audio. Moving forward, the development of VLMEvalKit will focus on expanding the repertoire of LMMs and benchmarks for video and other modalities. We are optimistic that this repository, along with all released resources, will contribute to advancing research in multi-modal learning.
@@ A.10 Miscellaneous
A.11 Video Understanding
  • MMBench-Video [31]: a long-form, multi-shot benchmark with about 600 YouTube videos (30 s–6 min) drawn from 16 everyday domains. It offers roughly 2000 volunteer-written Q&A pairs that probe 26 fine-grained skills and uses a GPT-4 adjudication pipeline for reliable scoring, making it a concise testbed for holistic video understanding.
  • Video-MME [34]: the first “full-spectrum” evaluation for video LLMs, containing 900 clips plus 2700 human-annotated Q&A pairs. Videos span short to hour-long segments across six visual domains, and the benchmark explicitly mixes frames, audio, and subtitles to examine cross-modal reasoning over varied temporal scales.
  • MVBench [68]: targets temporal reasoning by converting 20 classic vision tasks into dynamic versions that cannot be solved with a single frame. Automatic conversion of public annotations into multiple-choice QAs ensures fairness, and early results show mainstream MLLMs still fall short on many perception-to-cognition temporal skills.
  • MLVU [188]: focuses on long-video comprehension, gathering clips from 3 min up to 2 h and organising nine tasks that demand both global narrative understanding and fine-grained detail tracking. Initial evaluations (20 models, incl. GPT-4o) reveal that even the best system scores only 64.6%, underscoring the difficulty of sustained reasoning over extended contexts.
  • TempCompass [81]: an benchmark designed to gauge temporal perception. It covers diverse temporal aspects (action granularity, motion speed, event order, attribute change, etc.) and multiple task formats (multi-choice, yes/no, caption matching/generation). Carefully crafted “conflicting videos” reduce single-frame shortcuts, offering a stricter measure of whether models truly exploit temporal cues.
@@ B.1 Contributors with 3+ Major Contributions
  • PhoenixZ810 (Xiangyu Zhao): The contributor helped support WeMath [105], LogicVista [154], $M$-AlignBench [186], Video-ChatGPT [93], Chat-UniVI [57], and Llama-VID [71].
  • amitbcp (Amit Agarwal): The contributor helped support MUIRBench [131], Phi-3.5 [2], Idefics3 [60], VILA [75], xGen-$M$ [158], and MVTamperBench [4].
@@
Qualified Contributors (2025.06):
  • chenjix (JiXuan Chen): The contributor helped support Humanity’s last exam [102], LLama4 [127], ScreenSpot [22], ScreenSpot-v2 [152] and ScreenSpot-Pro [67].
  • Espere-1119-Song (Enxin Song): The contributor helped support Video-MMLU [114], MovieChat1k [113] and VDC [8].
  • maosong2022 (Song Mao): The contributor helped support MMVP [173], CVBench [126] and CharXiv [146], who also help solved many issues in the community.
  • SYuan03 (Shengyuan Ding): The contributor supports mPLUG-Owl3 [164], $M$-IFEval [27], ChartMimic [160] and fixed problems in Creation-MMBench [30] evaluation.
  • TianhaoLiang2000 (Tianhao Liang): The contributor supports mega-bench [11], Kimi-VL-A3B [124] and add vllm support for QwenVL/LLama4/InternVL Series Model.
  • zzc-1998 (Zicheng Zhang): The contributor supports Q-Bench [149], A-Bench [184] and Q-Bench Video [183].
@@ B.2 Full Contributor List
  • mary-0830: The contributor supports OmChat [185] and VLM-R1 [109].
@@
  • lihytotoro: The contributor supports MiniCPM-V-2.6 and MiniCPM-o-2.6 [163].
@@
  • FeipengMa6: The contributor supports WeMM [26] and WeThink-Qwen2.5VL-7B [161].
@@
  • lrlbbzl: The contributor of URSA-8B and URSA-8B-PS-GRPO [90].
  • lyccnb: The contributors of CG-Bench [9].
  • jscslld: The contributors of CG-Bench [9] and CG-AV-Counting [86].
@@
  • wulipc: The contributor of $C$-OCR [162].
  • mfarre: The contributor of SmolVLM and SmolVLM2.
@@
  • CaraJ7: The contributor of MathVerse [176] and MME-COT [56].
@@
  • bobo0810: The contributor of Phi-4 [1].
  • cmatachuan: The contributor of SAIL-VL-1.5 and SAIL-VL-1.6 [29].
  • dellixx: The contributor of Janus-Pro-1B [17].
  • g-h-chen: The contributor of UCSC-VLAA-Thinker [10].
  • maojialiang: The contributor of Ristretto.
  • RainJamesY: The contributor of VLM2Bench [175].
  • ryf1123: The contributor of VGRP-Bench [106].
  • suencgo: The contributor of Physics [32].
  • suyccc: The contributor of VMCBench [178].
  • sync-yxh: The contributor of Taichu-VLR.
  • tangkexian: The contributor of LEGO-Puzzle [119].
  • tianbinli: The contributor of MedXpertQA [191].
  • waltsun: The contributor of MOAT [166].
  • weiyao-wang: The contributor of AKI [143].
  • xingruiwang: The contributor of spatial457 [144].
  • xwy-bit: The contributor of VisuLogic [157].
  • yangtian6781: The contributor of MMCR [125].
  • zhaomh1998: The contributor of TDBench [49].
  • Hyggge: The contributor of valley2 [151].
  • xjtupanda: The contributor of $V^{*}$ Benchmark [150].
  • wutaiqiang: The contributor of PhyX [110].
  • mxin262: The contributor of OCR-Reasoning [52].
  • Zhouzone: The contributor of MSEarth [187] and OmniEarth [130].
  • An-LanWang: The contributor of WildDoc [129].
  • JunyingWang959: The contributor of A4Bench [134].
  • JiakangYuan: The contributor of MME-Reasoning [169].
  • donahowe: The contributor of Video-Holmes [21].
  • rohunagrawal: The contributor of Omni3D-Bench [94].
  • snowclipsed: The contributor of TallyQA Dataset [3].
  • JonasLoos: The contributor of SCAM Dataset [147].
  • zjh-tsinghua: The contributor of FlashVL [171].
  • nanocm: The contributor of OmniEarth [130] and XLRS-Bench-Lite [132].
  • zhouyiks: The contributor of MMVMBench
  • aarbelle: The contributor of GraniteVision [122].
  • yuzeng0-0: The contributor of VCR Bench [103].
  • Amber0614: The contributor of GOBench.